Definition 0.1. A sequence X = (xn) in R is said to converge to x E R, or x is said to be a limit of (xif for every e > 0 thePlease help! Only answer questions 5-8!
Answer #1

5) This statement is equivalent as for any real number L, there exists rational \epsilon'>0 such that EE , and then use the definition.

6) Note that this is not true consider the sequence 1). As then by this definition this sequence converges to both 1 and -1, which is a contradiction as the sequence does not converges by the original definition.

7) This is also not equivalent. Consider the same sequence. Then for all n\ge 1, (-1)-0 2 , but the sequence is not convergent .

8) That not equivalent. Note that this condition is a stronger condition (unlikely the other two) that is this statement implies the sequence is convergent but not conversely (by this definition only constant sequence is convergent). Consider the sequence \{\frac{1}{n}\} . Then this converges to 0. But note that choose \epsilon>\frac{1}{2} , then .이f€ , for n =.

Feel free to comment if you have any doubts. Cheers!